A pool table has a mass of 600 kg. In an attempt to move it, Peter exerts an upwards force of 1000 N on the pool table. What is the normal force now acting on the pool table?



1
Expert's answer
2021-07-12T12:16:24-0400

The total weight of the table (the force the table exerts on the floor) is:


"W = mg - 1000N"

where "m = 600kg" is the mass of the table, "g\approx 10N/kg" is the gravitational acceleration, and "1000N" is the Peter's force. The minus sign is due to the fact, that the gravity force "mg" is directed downward, and Peter's force is directed upward.

By the third Newton's law, the normal force acting on the pool table is equal to its weight:


"N = W = 600\cdot 10 - 1000 = 5000N"

Answer. 5000N.